# Meta-analysis of median survival times with inverse-variance weighting

### v1

@@ 4 Application
In this section, we illustrate an application of the Wald approximation-based approach to a recent meta-analysis which investigated differences in survival between individuals randomized to experimental groups versus comparator groups in randomized trials for cancer therapies [9]. For the purposes of this illustration, we focus on randomized trials in non-small cell lung cancer (NSCLC) and consider the outcome of overall survival (OS).
A total of 30 studies reported estimates of median OS and 95% confidence intervals separately in the experimental and comparator groups (see Appendix D). These studies had an average sample size of 228 in the experimental arm and 205 in the comparator arm. Differences of median OS ranged from -5.39 months (favoring comparator group) to 12.40 months (favoring experimental group). Ratios of median OS ranged from 0.57 (favoring comparator group) to 1.82 (favoring experimental group).
@@
The pooled estimate of median survival in the comparator group was 12.81 months (95% CI: 10.85, 14.77). As one may expect, there was a substantial amount of between-study heterogeneity. The 95% prediction interval was wide (2.85 to 22.77), and the estimated index was 95.03%.
@@
Recall that the meta-analysis of the difference of medians and the meta-analysis of the ratio of medians methods involve different assumptions on the form of the between-study heterogeneity. That is, the difference of medians meta-analysis assumes heterogeneity on the additive scale (i.e., the true study-specific differences of medians are normally distributed) and the ratio of medians meta-analysis assumes heterogeneity on the multiplicative scale (i.e., the true study-specific log ratios of medians are normally distributed). Consequently, when applying both of these methods to the same data, such as in this illustration, one may be concerned about misspecifying the form of between-study heterogeneity in at least one of these methods. In Appendix C.2, we evaluate the performance of these two methods in our simulation study when misspecifying the form of the between-study heterogeneity in this manner. We found that such misspecification had little impact on the performance of these methods. Related, see [32] for a discussion of implications of the distributional assumptions on between-study heterogeneity.
@@ 5 Discussion
In this article, we present an inverse-variance weighted approach to meta-analyzing median survival times when primary studies report Kaplan-Meier estimates of median survival times and confidence intervals. In particular, we described methods to meta-analyze the median survival time, the difference of median survival times between two groups, and the ratio of median survival times between two groups. A key aspect of these methods is that they estimate the standard errors of the study-specific median survival estimates by treating the reported confidence intervals as Wald intervals. Our simulation results suggest that this may often be a reasonable approximation in practice for confidence interval methods that are commonly used in this context. The Wald approximation-based approach generally did not systematically over- or under-estimate the standard error of median survival regardless of the method used to construct the confidence interval. These results held for several different event time distributions, censoring time distributions, and sample sizes. The Wald approximation-based approach also performed well in our meta-analysis simulations. In particular, this approach was comparable to a benchmark approach that uses the actual true within-study standard errors.
@@ 5.1 Related methods
Flores and Müller [39] proposed a Bayesian nonparametric approach to meta-analyze outcome measures based on median survival in the context of cancer immunotherapy studies. This approach assumes that the confidence intervals around the medians are constructed based on central order statistics, which can be seen as an analogous assumption as that in the Wald approximation (see Remark 1). Poli et al. [40] proposed an alternative Bayesian nonparametric approach for meta-analyzing survival times in this context that involves a different construction for the priors. A key strength of these approaches is that they accommodate flexible patterns of heterogeneity across studies. These approaches may be an excellent option for meta-analysts who are familiar with Bayesian nonparametrics and can tailor the estimation algorithm to their application at hand.
@@ 5.2 Limitations and future work
Our simulation results also suggest that caution should be taken when applying the Wald approximation-based approach with small study-specific sample sizes (e.g., ). The (relative) bias and variability of the standard error estimator was largest in these settings. This is exacerbated when the event time distribution is highly skewed. In Appendix E, we perform additional simulations with extremely skewed event time distributions to illustrate this point. However, small sample sizes and highly skewed event time distributions would likely pose challenges for alternative methods as well.
Several modifications to the methods described in this work can be considered. In settings where the Wald approximation may poorly estimate the within-study standard errors, one may consider using a jackknife approach (e.g., see Ozturk and Balakrishnan [17]), nonparametric bootstrap, or robust variance estimation [50, 51] to construct a confidence interval around the pooled estimate. Such approaches are less sensitive to the study-specific weights and may yield better confidence interval coverage in these settings. We did not adopt such approaches here because the Wald approximation performed well for estimating the within-study standard errors. Another modification that may be considered is meta-analyzing the median survival time based on a log transformation rather than on the original scale, as used for the ratio of medians meta-analysis. A key implication of meta-analyzing the untransformed or log transformed median is whether heterogeneity of the study-specific median survival times is thought to be additive or multiplicative. Based on our simulations in Appendix C.2, we suspect that this distinction would not strongly affect the performance of the method.
@@ Acknowledgements
We would like to thank Renata Iskander for helping collect the data set used in the example. The simulations in this work were run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University.
@@
The authors declare no potential conflict of interest.
Data and code availability
@@
The data and code used in the simulation study and data application are available at
https://github.com/stmcg/ma-median-survival.
Appendix A Review of Kaplan-Meier estimation
In this section, we briefly review aspects of Kaplan-Meier estimation that are used throughout the main text, following notational conventions in Andersen [16]. Proofs of all results can be found in Andersen [16].
A.1 Estimation of the survival function
Let denote the cumulative distribution function and denote the probability density function of the event times. Let denote the survival function. Throughout this section, we let where . Let denote the observed event times.
The Kaplan-Meier estimator of is given by
where denotes the number of events at time and denotes the number of individuals at risk prior to time . Under certain regularity conditions, is asymptotically normal with mean and variance that can be estimated by the Greenwood formula:
One can construct Wald-type confidence intervals for based on these estimators. For example, a confidence interval is given by
where denotes the quantile of the standard normal distribution. Since , one often applies a transformation (e.g., log or log-minus-log) to construct the confidence interval for better small sample properties. A confidence interval based on the log transformation is given by
and a confidence interval based on the log-minus-log transformation is given by
A.2 Estimation of quantiles
We next consider estimation of quantiles of the distribution of event times. The th quantile and the Kaplan-Meier estimator of the th quantile are defined by
Under certain regularity conditions, is asymptotically normal with mean and variance that can be estimated by
The estimator may be based on a kernel function estimator of the density , for example.
One can construct a Wald-type confidence interval for by
Since , one may consider applying a transformation such as the log transformation for better small sample properties. However, a disadvantage for Wald-type confidence intervals for is that they require an estimate of the probability density function .
A more commonly used approach to construct a confidence interval for is the Brookmeyer-Crowley method. This approach inverts a hypothesis test of the null versus the alternative based on the asymptotic normality of (or for some transformation ). That is, the interval is the set of all such that one fails to reject the null hypothesis when testing versus at the significance level. More formally, the confidence interval is given by
An advantage of this approach is that it does not require estimating .
Appendix B Impact of asymmetry in confidence interval tail probabilities
In this section, we provide some intuition on the impact of asymmetry in confidence interval tail probabilities (i.e., and ) on the Wald approximation in (3).
To fix ideas, suppose now that the confidence interval is of the form
where are unknown constants that depend on . Assuming that the standard error estimator is consistent, it is straightforward to see that in large samples and where and such that . Therefore, if and are were known, we could back-compute the standard error estimate similar to (3), i.e.,
Analogous to Remark 1 in the main text, one can alternatively motivate the standard error estimator in (4) when and are quantile estimators with and .
Although the values of and may be unknown to the meta-analyst, we can compare (i.e., appearing in (4)) and (i.e., appearing in (3)) across a range of and values. One can see that reaches a maximum of when . Furthermore, unless is nearly equal to 0 or (and thus is nearly or 0), is close to its maximum of . We illustrate this in Figure 6 and Table 4 for , noting that similar conclusions hold for other common levels.
In summary, when confidence intervals are asymmetric in their tail probabilities, the Wald approximation may overestimate the standard error of median survival and the bias may increase as the degree of asymmetry in the confidence interval increases. However, the magnitude of the bias may often be small unless one of the tail probabilities is nearly 0.
| 0.025 | 0.025 | 1.96 | 1.96 | 0.26 | 0.26 |
| 0.030 | 0.020 | 1.88 | 2.05 | 0.25 | 0.26 |
| 0.035 | 0.015 | 1.81 | 2.17 | 0.25 | 0.26 |
| 0.040 | 0.010 | 1.75 | 2.33 | 0.25 | 0.26 |
| 0.045 | 0.005 | 1.70 | 2.58 | 0.23 | 0.26 |
| 0.049 | 0.001 | 1.65 | 3.09 | 0.21 | 0.26 |
Appendix C Additional simulation results
C.1 Study-level simulations
| Censoring Time Distribution | ||
| Event Time Distribution | Uniform | Exponential |
| Exponential | 37 | 41 |
| Weibull | 31 | 38 |
| Weibull Mixture | 26 | 33 |
| Sample Size | ||||||
| Event Time Dist. | Censoring Dist. | CI Method | ||||
| Exponential | Uniform | BC (log) | 14 | 8 | 4 | 0 |
| BC (log-log) | 4 | 2 | 1 | -1 | ||
| Bootstrap | 4 | 1 | 1 | -1 | ||
| Exponential | BC (log) | 12 | 10 | 3 | 1 | |
| BC (log-log) | 4 | 1 | 1 | 0 | ||
| Bootstrap | 3 | 1 | 0 | 0 | ||
| Weibull | Uniform | BC (log) | 11 | 3 | 0 | 0 |
| BC (log-log) | 4 | 0 | -1 | 0 | ||
| Bootstrap | 1 | -2 | -2 | -1 | ||
| Exponential | BC (log) | 12 | 5 | 2 | 1 | |
| BC (log-log) | 2 | 1 | 1 | 1 | ||
| Bootstrap | 0 | -1 | 0 | 1 | ||
| Weibull Mixture | Uniform | BC (log) | 16 | 6 | 2 | 0 |
| BC (log-log) | 4 | 2 | 0 | 0 | ||
| Bootstrap | 2 | 1 | -1 | 0 | ||
| Exponential | BC (log) | 19 | 6 | 2 | 2 | |
| BC (log-log) | 4 | 1 | 0 | 1 | ||
| Bootstrap | 2 | 0 | 0 | 1 | ||
C.2 Meta-analytic level simulations: misspecifying the form of between-study heterogeneity
In this subsection, we conduct simulations that evaluate the performance of the Wald approximation-based approach when misspecifying the form of the between-study heterogeneity. One motivation for these simulations arises from the fact that the method for meta-analyzing the difference of medians involves different distributional assumptions on the between-study heterogeneity compared to the method for meta-analyzing the ratio of medians, yet researchers may wish to apply both methods to the same data (e.g., see Section 4 in the main text). These simulations involve meta-analyzing the difference of medians in scenarios where the heterogeneity assumptions for meta-analyzing the ratio of medians (rather than the difference of medians) actually hold. Additionally, these simulations involve meta-analyzing the ratio of medians in scenarios where the heterogeneity assumptions for meta-analyzing the difference of medians hold.
C.2.1 Meta-analysis of the difference of median survival
Data generation
We used the data generating mechanism described in Section 3.2.3 of the main text. In these settings, recall that the true study-specific log ratios of medians are normally distributed across studies. The distributions of the true study-specific differences of medians in these settings are illustrated in Figure 9, which were obtained by Monte Carlo integration with samples.
Note that the degree of between-study variance was higher in these simulations compared to those in Section 3.2.2 of the main text. The variance of the true study-specific difference of medians was approximately 24.14 in the high heterogeneity setting and 7.80 in the low heterogeneity setting here. For comparison, recall that the variance of the true study-specific differences of medians was 12 in the high heterogeneity setting and 4 in the low heterogeneity setting in the simulations in Section 3.2.2.
We applied the Wald approximation-based approach to meta-analyze the difference of medians. We also applied the benchmark approach which uses the true standard errors of the study-specific differences of medians, as described in the main text. Like the Wald approximation-based approach, note that the benchmark approach also misspecifies the distribution study-specific difference of medians in these simulations. That is, the benchmark approach represents the best-case scenario of the Wald approximation-based approach with respect to estimating the within-study standard errors.
Results
The results are summarized in Figure 10 and Table 7. The trends are similar to those of the simulation study in Section 3.2.2 of the main text. The Wald approximation-based approach performed similarly to the benchmark approach in each setting. As one would expect, the standard errors of the estimators were generally larger here compared to those in Section 3.2.2 due to the higher degree of between-study heterogeneity.
| Wald Approx. Approach | Benchmark Approach | ||||||
| Pooled Dif. of Medians | High | -0.26 | 1.33 | 0.93 | -0.25 | 1.34 | 0.93 |
| Moderate | -0.17 | 0.92 | 0.94 | -0.16 | 0.93 | 0.94 | |
| None | -0.02 | 0.63 | 0.96 | -0.02 | 0.63 | 0.96 | |
| Between-Study Variance | High | -2.10 | 10.73 | 0.95 | -1.51 | 10.98 | 0.95 |
| Moderate | -0.43 | 5.08 | 0.96 | -0.18 | 5.22 | 0.95 | |
C.2.2 Meta-analysis of the ratio of median survival
Data generation
We used the data generating mechanism described in Section 3.2.2 of the main text. Here, the true study-specific differences of medians are normally distributed across studies. The distributions of the true study-specific log ratios of medians in these settings are illustrated in Figure 11, which were obtained by Monte Carlo integration with samples.
The degree of between-study variance is lower in these simulations compared to those in Section 3.2.3 of the main text. The variance of the true study-specific log ratio of medians is approximately 0.016 in the high heterogeneity setting and 0.005 in the low heterogeneity setting here. Recall that the variance of the true study-specific log ratio of medians was 0.03 in the high heterogeneity setting and 0.01 in the low heterogeneity settings in the simulations in Section 3.2.3.
We applied the Wald approximation-based approach and benchmark approach to meta-analyze the ratio of medians.
Results
The results are summarized in Figure 12 and Table 8. We observed similar trends as those in Section 3.2.3 of the main text. The Wald approximation-based approach and benchmark approach performed similarly. As expected, the standard errors of the estimators were generally smaller here compared to those in Section 3.2.2 due to the lower degree of between-study heterogeneity.
| Wald Approx. Approach | Benchmark Approach | ||||||
| Pooled Ratio of Medians | High | -0.73 | 3.61 | 0.96 | -0.73 | 3.58 | 0.97 |
| Moderate | -0.34 | 2.86 | 0.95 | -0.35 | 2.82 | 0.96 | |
| None | 0.01 | 2.20 | 0.97 | 0.02 | 2.16 | 0.97 | |
| Between-Study Variance | High | 0.07 | 0.98 | 0.94 | 0.08 | 0.96 | 0.94 |
| Moderate | 0.05 | 0.48 | 0.96 | 0.05 | 0.48 | 0.96 | |
Appendix D Additional results for the data application
| Experimental Group | Comparator Group | |||
| Study | Median OS (95% CI) | Median OS (95% CI) | ||
| NCT00946712 | 656 | 10.90 (9.50, 12.00) | 657 | 9.20 (8.70, 10.30) |
| NCT01041781 | 154 | 11.40 (9.46, 14.06) | 158 | 12.50 (9.36, 13.83) |
| NCT01386385 | 18 | 27.60 (17.40, 27.60) | 13 | 15.20 (6.60, 20.60) |
| NCT01395758 | 51 | 6.80 (4.97, 10.70) | 45 | 8.50 (6.37, 13.97) |
| NCT01466660 | 160 | 27.86 (25.13, 32.85) | 159 | 24.54 (20.57, 28.88) |
| NCT01642251 | 64 | 10.30 (8.90, 12.00) | 64 | 8.90 (8.30, 11.30) |
| NCT01828112 | 115 | 18.10 (13.40, 23.90) | 116 | 20.10 (11.90, 25.10) |
| NCT01933932 | 254 | 8.70 (3.60, 16.80) | 256 | 7.90 (3.80, 20.10) |
| NCT01951586 | 148 | 10.70 (8.54, 12.35) | 78 | 10.90 (9.26, 15.54) |
| NCT02008227 | 425 | 13.80 (11.80, 15.70) | 425 | 9.60 (8.60, 11.20) |
| NCT02041533 | 211 | 14.36 (11.60, 17.45) | 212 | 13.21 (10.68, 17.08) |
| NCT02151981 | 279 | 26.80 (23.50, 31.50) | 140 | 22.50 (20.20, 28.80) |
| NCT02152631 | 270 | 7.40 (6.50, 8.80) | 183 | 7.80 (6.40, 9.50) |
| NCT02264990 | 298 | 12.10 (10.40, 14.90) | 297 | 12.10 (10.00, 13.70) |
| NCT02296125 | 279 | 38.60 (34.50, 41.80) | 277 | 31.80 (26.60, 36.00) |
| NCT02352948 | 62 | 11.70 (8.20, 17.40) | 64 | 6.80 (4.90, 10.20) |
| NCT02352948 | 174 | 11.50 (8.70, 14.10) | 118 | 8.70 (6.50, 11.70) |
| NCT02366143 | 359 | 19.20 (17.00, 23.80) | 337 | 14.70 (13.30, 16.90) |
| NCT02367781 | 456 | 18.60 (15.80, 21.20) | 229 | 13.90 (12.00, 18.70) |
| NCT02367794 | 343 | 14.20 (12.30, 16.80) | 340 | 13.50 (12.20, 15.10) |
| NCT02387216 | 71 | 7.70 (3.60, 10.40) | 38 | 8.40 (5.80, 14.70) |
| NCT02395172 | 396 | 10.60 (9.20, 12.30) | 396 | 9.90 (8.10, 11.90) |
| NCT02409342 | 277 | 18.90 (13.40, 23.00) | 277 | 14.70 (11.20, 16.50) |
| NCT02450539 | 106 | 7.00 (5.00, 8.78) | 53 | 12.39 (7.13, 15.98) |
| NCT02453282 | 163 | 11.90 (9.00, 17.70) | 162 | 12.90 (10.50, 15.00) |
| NCT02453282 | 163 | 16.30 (12.20, 20.80) | 162 | 12.90 (10.50, 15.00) |
| NCT02542293 | 410 | 10.90 (9.30, 12.60) | 413 | 12.10 (10.30, 13.50) |
| NCT02657434 | 292 | 17.50 (13.20, 19.60) | 286 | 13.60 (11.00, 15.70) |
| NCT02785952 | 125 | 10.00 (8.00, 14.40) | 127 | 11.00 (8.60, 13.70) |
| NCT02855125 | 64 | 7.92 (6.28, 10.78) | 64 | 9.82 (7.66, 13.40) |
Appendix E Simulations with extreme skewness
In this section, we illustrate the performance of the Wald approximation under settings with increasingly high skewness in the event time distribution.
E.1 Data generation
We adopt the same simulation design as in the study-level simulations in the main text (i.e., Section 3.1) with the following changes. Here, we generated the underlying event times from either a Weibull distribution with or . Figure 13 illustrates the probability density functions of the two distributions. We used the distributions for censoring and used the same four sample sizes as in the main text.
Note that decreasing the shape parameter in this context increases the skewness of the event time distribution, which can be seen in the following ways. Pearson’s moment coefficient of skewness is approximately 3.8 for the distribution and is 19.6 for the distribution. An alternative measure of skewness called Bowley’s coefficient of skewness [53] has been commonly used in the context of meta-analyzing median-based outcome measures (e.g., see [47, 48, 17, 11, 12]). Bowley’s skewness ranges from to , where negative values indicate left skewness and positive values indicate right skewness. Bowley’s skewness of the and distributions are approximately 0.43 and 0.77, respectively.
When the underlying event times followed the Weibull distribution with , the percentage of censoring was 34% in the scenarios with the uniform censoring time distribution and 38% in the scenarios with the exponential censoring time distribution. When the underlying event times followed the Weibull distribution with , the percentage of censoring was 36% in the scenarios with the uniform censoring time distribution and 38% in the scenarios with the exponential censoring time distribution.
E.2 Results
The simulation results are presented in Figure 14 and Table 10. The relative bias of the standard error estimators in scenarios with the event time distribution were comparable to that observed in the simulations in the main text. However, for the more highly skewed event time distribution, we observed larger relative bias values especially when the Brookmeyer-Crowley method with the log transformation was used to construct confidence intervals. For example, in the scenario with exponential censoring and a sample size of 100, the relative bias reached -19%. Consistent with the simulation results from the main text, the relative bias decreased as the sample size increased. When the sample size was 1000, the relative bias was below 2% across all scenarios.
| Sample Size | ||||||
| Event Time Dist. | Censoring Dist. | CI Method | ||||
| Uniform | BC (log) | 10 | 12 | 5 | 1 | |
| BC (log-log) | 3 | 3 | 2 | 0 | ||
| Bootstrap | 3 | 3 | 2 | 0 | ||
| Exponential | BC (log) | 12 | 16 | 6 | 0 | |
| BC (log-log) | 2 | 4 | 2 | -1 | ||
| Bootstrap | 2 | 3 | 2 | -1 | ||
| Uniform | BC (log) | -12 | 16 | 9 | 1 | |
| BC (log-log) | -7 | 5 | 3 | 0 | ||
| Bootstrap | -7 | 6 | 4 | 0 | ||
| Exponential | BC (log) | -15 | 19 | 9 | 2 | |
| BC (log-log) | -8 | 5 | 3 | 1 | ||
| Bootstrap | -4 | 8 | 3 | 0 | ||

### v2

@@ 4 Application
In this section, we illustrate an application of the Wald approximation-based approach to a recent meta-analysis which investigated differences in survival between individuals randomized to experimental groups versus comparator groups in randomized trials for cancer therapies [9]. For the purposes of this illustration, we focus on randomized trials in non-small cell lung cancer (NSCLC) and consider the outcome of overall survival (OS). The data and code used for this application are available at https://github.com/stmcg/ma-median-survival.
A total of 28 studies reported estimates of median OS and 95% confidence intervals separately in the experimental and comparator groups (see Section 6 of the Supplementary Material). The average sample size in the experimental group was 228 and the average sample size in the comparator group was 205. Differences of median OS ranged from -5.39 months (favoring comparator group) to 12.40 months (favoring experimental group). Ratios of median OS ranged from 0.57 (favoring comparator group) to 1.82 (favoring experimental group).
@@
The pooled median survival estimate in the comparator group was 12.81 months (95% CI: 10.85, 14.77). As one may expect, there was a substantial amount of between-study heterogeneity. The 95% prediction interval was wide (2.85 to 22.77), and the estimated index was 95.03%.
@@
Two of the 28 studies (NCT02352948 and NCT02453282) each contributed two sets of estimates to our meta-analysis. NCT02352948 consisted of two independent sub-studies, and NCT02453282 included two independent experimental groups. Similar results were obtained when using multivariate meta-analytic approaches that account for potential correlation arising from these two studies providing multiple estimates (e.g., multivariate models with study-specific random intercepts). For simplicity and consistency with the rest of the paper, we presented the univariate analysis here.
@@ 5 Discussion
In this article, we present an inverse-variance weighted approach to meta-analyzing median survival times when primary studies report Kaplan-Meier estimates of median survival times and confidence intervals. In particular, we described methods to meta-analyze the median survival time, the difference of median survival times between two groups, and the ratio of median survival times between two groups. A key aspect of these methods is that they estimate the standard errors of the study-specific median survival estimates by treating the reported confidence intervals as Wald intervals.
Our simulation results suggest that this approach may often perform well in practice for commonly used confidence interval methods (i.e., the Brookmeyer-Crowley method and nonparametric bootstrap) when the effective sample size is reasonably large. The Wald approximation-based approach generally did not systematically over- or under-estimate the standard error of median survival in such scenarios regardless of the method used to construct the confidence interval. These results held for several different event time distributions, censoring time distributions, and sample sizes. The Wald approximation-based approach also performed well in the corresponding meta-analysis simulations. This approach was approximately unbiased for estimating the pooled outcome measure and between-study variance and had corresponding near-nominal confidence interval coverage in each setting. Furthermore, this approach was comparable to a benchmark approach that uses the actual true within-study standard errors.
In settings where the effective sample size is small, such as due to a small number of participants and a high degree of censoring, the Wald approximation can perform poorly for estimating within-study standard errors. This can in turn lead to biased estimation of between-study heterogeneity in random effects meta-analyses and poor coverage of the pooled outcome measure in common effect meta-analyses.
@@ 5.1 Related methods
Flores and Müller [37] proposed a Bayesian nonparametric approach to meta-analyze outcome measures based on median survival in the context of cancer immunotherapy studies. This approach assumes that the confidence intervals around the medians are constructed based on central order statistics, which can be seen as an analogous assumption as that in the Wald approximation. Poli et al. [38] proposed an alternative Bayesian nonparametric approach for meta-analyzing survival times in this context that involves a different construction for the priors. A key strength of these approaches is that they accommodate flexible patterns of heterogeneity across studies. These approaches may be an excellent option for meta-analysts who are familiar with Bayesian nonparametrics and can tailor the estimation algorithm to their application at hand.
@@ 5.2 Limitations and future work
The simulation studies focused on traditional parametric event time and censoring distributions. Results may differ under more realistically complex distributions, which may be interesting to explore in future work. Moreover, to keep the results digestible, our meta-analysis simulations considered relatively limited settings (i.e., a total of 18 scenarios based on varying the outcome type, heterogeneity level, and sample sizes of the primary studies). Future work could explore the impact of a number of other factors, such as the number of primary studies, event time and censoring distributions, and random effects distribution. In our analyses, we found that the effective sample sizes of the primary studies was the strongest factor affecting the performance of the Wald approximation.
Several modifications to the methods described in this work can be considered. In settings where the Wald approximation may poorly estimate the within-study standard errors, one may consider using a jackknife approach (e.g., see Ozturk and Balakrishnan [17]), nonparametric bootstrap, or robust variance estimation [49, 50] to construct a confidence interval around the pooled estimate. Such approaches are less sensitive to the study-specific weights and may yield better confidence interval coverage in these settings. These approaches may be particularly useful in settings where primary studies have small effective sample sizes, as the Wald approximation can perform poorly in such settings. Another modification that may be considered is meta-analyzing the median survival time based on a log transformation rather than on the original scale, as used for the ratio of medians meta-analysis. A key implication of meta-analyzing the untransformed or log transformed median is whether heterogeneity of the study-specific median survival times is thought to be additive or multiplicative. Based on our simulations in Section 5.1 of the Supplementary Material, we suspect that this distinction would not strongly affect the performance of the method.
@@ Acknowledgements
We would like to thank Renata Iskander for helping collect the data set used in the example. This work was primarily conducted while SM was at Harvard University. The simulations in this work were run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University.
Financial disclosure
The authors declare that no specific funding has been received for this article.
@@
The authors declare no potential conflict of interests.
Open Research
@@
The data and code used in the simulation study and data application are available at https://github.com/stmcg/ma-median-survival.